OpenStudy (anonymous):

where does the graph of y=-2x-16 cross the x-axis?

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x y=-2x-16 0=-2x-16 0+16 = -2x-16+16 16 = -2x ... ... ... x = ???

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

couldn't you just graph it?